2 C2H6(g) +7 O2(g) = 4 CO2(g) + 6 H2O(g)

1. Calculate the volume in liters of CO2 at STP, that could be produced from 7.25 g C2H6.

2. Calculate the volume of O2 at 100oC and 1.25 atm, that would react with 7.52g of C2H6

1
Expert's answer
2021-10-14T18:39:02-0400

1. M(C2H6) = 30.07 g/mol

n(C2H6) "= \frac{7.25}{30.07} = 0.241 \;mol"

According to the reaction equation

n(CO2) = 2n(C2H6) "= 2 \times 0.241 = 0.482 \;mol"

At STP one mole of any gas occupies 22.4 L.

V(CO2) "= 22.4 \times 0.482 = 10.80 \;L"

2. n(C2H6) "= \frac{7.52}{30.07} = 0.25 \;mol"

According to the reaction equation

n(O2) "= \frac{7}{2}" n(C2H6) "= \frac{7}{2} \times 0.25 = 0.875 \;mol"

Ideal gas law:

"pV=nRT \\

p = 1.25 \;atm \\

T = 100 +273 = 373 \;K"

R = 0.08206 L×atm/mol×K

"V(O_2) = \frac{nRT}{p} \\

= \frac{0.875 \times 0.08206 \times 373}{1.25} \\

= 21.42 \;L"